What is the age in years of a bone in which the 14C/12C ratio is measured to be 3.9×10−13?
Answer #1

t = -(t1/2/ln(2)) * ln(xt/x0)

t1/2 = half life of 14-C = 5730 yrs.

xt/x0 = 3.9*10^(-13)


t = -(5730/ln2)*ln(3.9*10^(-13))

   = 236199.7 years